What mass of calcium carbonate require to produce 400ml of co2 on heating
1
Expert's answer
2019-06-03T05:48:36-0400

CaCO3 --> CaO + CO2

According to ideal gas law:

PV=nRT

Vm=n/V=RT/P

Volume of ideal gas at STP is equal to 22.4L for 1 mol.

Vm=22.4 L/mol

n(CO2)=V(CO2)/Vm

n(CO2)=0.4/22.4=0.018mol

n(CO2):n(CaCO3)=1:1

n(CaCO3)=0.018 mol

m(CaCO3)=n(CaCO3)*M(CaCO3)

m(CaCO3)=0.018*100=1.8g


Answer: m(CaCO3)=1.8g